OpenStudy (anonymous):

Solve for solutions in interval [0 degrees, 360 degrees). Give solutions to the nearest tenth as appropriate. cot (θ)+2csc (θ)=3

OpenStudy (dumbcow):

\[\rightarrow \frac{\cos \theta}{\sin \theta} + \frac{2}{\sin \theta} = 3\] \[\cos \theta +2 = 3 \sin \theta\] square both sides \[\cos^2 \theta +4\cos \theta +4 = 9\sin^2 \theta = 9(1-\cos^2 \theta)\] \[10 \cos^2 \theta +4 \cos \theta -5 = 0\] solve using quadratic formula

OpenStudy (anonymous):

Thanks. After using the quadratic equation the answer it gave me was the following  When I input the values I get small values. How can I have my answer in degrees?